# Taming Data Challenges in ML-based Security Tasks: Lessons from Integrating Generative AI

### v1

@@ Abstract
Machine learning-based supervised classifiers are widely used for security tasks, and their improvement has been largely focused on algorithmic advancements. We argue that data challenges that negatively impact the performance of these classifiers have received limited attention. We address the following research question: Can developments in Generative AI (GenAI) address these data challenges and improve classifier performance? We propose augmenting training datasets with synthetic data generated using GenAI techniques to improve classifier generalization. We evaluate this approach across 7 diverse security tasks using 6 state-of-the-art GenAI methods and introduce a novel GenAI scheme called Nimai that enables highly controlled data synthesis. We find that GenAI techniques can significantly improve the performance of security classifiers, achieving improvements of up to 32.6% even in severely data-constrained settings (only $\sim$180 training samples). Furthermore, we demonstrate that GenAI can facilitate rapid adaptation to concept drift post-deployment, requiring minimal labeling in the adjustment process. Despite successes, our study finds that some GenAI schemes struggle to initialize (train and produce data) on certain security tasks. We also identify characteristics of specific tasks, such as noisy labels, overlapping class distributions, and sparse feature vectors, which hinder performance boost using GenAI. We believe that our study will drive the development of future GenAI tools designed for security tasks.
I Introduction
@@
The application of machine learning (ML) to counteract cybersecurity threats is becoming increasingly prevalent. A widely adopted approach involves creating ML-based security classifiers through supervised learning to identify or categorize threats. These classifiers serve to identify or categorize malicious users or behaviors on online platforms [108], malicious software [110], network breaches [71], harmful network traffic/devices/entities [96, 88], software vulnerabilities [118], and web security hazards [112]. These defense mechanisms emphasize lowering error rates (false alarms), enhancing generalization to unfamiliar test scenarios, and adding robustness against adaptive attackers. In pursuing these goals, primary attention has been paid to addressing the algorithmic challenges, which pertain to developing effective ML models and feature engineering techniques. However, the data challenges that affect the practical effectiveness of these defenses have received limited attention, resulting in significant limitations in further advancing performance. Our examination uncovers the following crucial data challenges in security (Section II): significant class imbalance, insufficient representation of attack patterns, inadequate training samples, high-dimensional features, and concept drift [111].
@@
In this study, we explore how recent advancements in AI can alleviate data-related challenges to enhance classification performance. This leads to our primary research inquiry: What are the challenges and opportunities associated with employing Generative Artificial Intelligence (GenAI) to address data challenges in machine learning (ML)-based security applications? GenAI models are capable of learning the distribution of a dataset and generating varied synthetic instances [83]. A pivotal concept is utilizing carefully crafted synthetic data from GenAI to augment the existing training dataset of a security classifier, thereby enhancing generalization capabilities. Figure 1 (a) depicts the process of leveraging GenAI for data augmentation to refine performance before the classifier’s deployment. We specifically concentrate on tabular data tasks, as this represents a prevalent methodology for developing ML-based security classifiers. Examples of such tasks are presented in Sections II and IV.
Employing GenAI to tackle data challenges in security is difficult for multiple reasons:
@@
  • Data synthesis in the security domain is inherently harder. Unlike other domains, e.g., computer vision [67, 69, 22], data in the security domain is inherently “adversarial” and captures behavior of an entity aiming to evade defenses. Thus, defenders typically have less knowledge of attack data distributions than of benign data. A GenAI scheme has to work with biased and limited attack data, as well as data with noisy samples (erroneous labels).
@@
  • Existing state-of-the-art GenAI models for tabular data are not tailor-made for security tasks. Several GenAI models for tabular data have emerged recently, e.g., TVAE [106], CTAB-GAN+ [117], TabDDPM [57], TabSyn [115], REaLTabFormer [92] and GReaT [16]. No study has systematically tested these GenAI schemes on security tasks. It is unclear how these GenAI schemes (for tabular data) can adapt to the data challenges in security.
@@
  • A GenAI scheme for security should not only address data challenges before deployment, but also after deployment. Security classifiers are known to suffer from concept drift due to adaptive attackers [6]. It is unclear how GenAI schemes can help with recovery from concept drift with minimal data labeling effort.
Figure 1: Overview of data augmentation using GenAI.
@@
  • We start by highlighting key data challenges impacting development of security classifiers. We review 35 papers published in top security venues. We find 32/35 papers report at least one data challenge negatively impacting their classifier performance. We use our findings to select 7 security tasks suffering from these data challenges for evaluation. This includes diverse tasks: malware classification [110, 5, 96], OS fingerprinting [48], BGP hijacking detection [97], Tor-based malware traffic detection [31], and web cookie privacy classification [15].
  • We present the first systematic study of the effectiveness of SoTA GenAI schemes (designed for tabular data) to address data challenges and boost performance of security classifiers. We evaluate 6 SoTA GenAI schemes on our 7 security tasks. This includes GenAI models based on VAE, LLM, GAN and Diffusion models. We use synthetic data from GenAI schemes to augment training data for security tasks and boost performance.
  • We identify a missing capability in existing GenAI schemes (for tabular data)—none provide any control to target specific regions of the data manifold to alleviate complex biases. To fill this gap, we propose Nimai, a novel VAE-based scheme, which uses a discrete latent space to enable controlled sample-conditioned generations. Given an existing sample, Nimai can generate samples in its vicinity, providing more fine-grained control to the defender to address bias issues. Our analysis shows that carefully crafted synthetic data improves classifier generalization by mitigating biases in the data.
@@
    Opportunities before deployment: This part focuses on mitigating data challenges before deployment of the classifier. We measure the performance gain (or degradation) from data augmentation compared to using only real training data. We obtain two key findings: (1) Existing GenAI schemes, among TVAE, CTAB-GAN+ and TabDDPM demonstrate performance improvement in 2 to 4 tasks, e.g., with up to 14.8% gain in the BGP hijacking detection task. (2) Our new scheme, Nimai which offers highly controlled data generation, achieved the best performance gain in 3/4 tasks.
    In an extremely data-challenged BGP hijacking detection task ($<200$ training samples), Nimai outperformed all baselines with up to a 32.6% gain. Our findings demonstrates the significant untapped potential of GenAI in security.
@@
    Challenges: Our analysis also revealed significant challenges with using GenAI: (1) We discover fundamental challenges with instantiating (training and generating data) existing GenAI models on our diverse tasks. 5/6 existing GenAI schemes failed to instantiate in at least one of the 7 tasks. For example, the LLM-based GenAI scheme called GReaT [16] failed to generate data in 6/7 tasks due to scalability issues. (2) We identify 3 tasks with shared challenges of limited or biased data, and added complexities for GenAI: noisy labels, sparse features, and class overlap. Most GenAI schemes, including our new approach, fail to improve performance, highlighting key directions for future work.
We anticipate that our findings will lead to the creation of a GenAI tool specifically designed for security classification tasks, effectively tackling all the challenges we highlighted. In Section IX, we outline several specific directions for future work. We will make all our code and datasets accessible to the community.
II Data Challenges in Security Defenses
We conduct a study to characterize data challenges faced by researchers in ML-based cybersecurity tasks. We review 35 papers  [^1] Listed in Appendix Table VIII. published in leading venues (e.g., IEEE S&P, USENIX Security and IMC) that developed at least one security classifier. The articles are semi-automatically selected based on their relevance to ML in security research. [^2] Our paper selection methodology is detailed in Appendix Section -B.
We find that 32/35 articles explicitly reported at least one data challenge that negatively affected classifier performance. We identify the following broad data challenges:
@@ II Data Challenges in Security Defenses
a) Imbalanced class distribution (10/35 papers). Certain classes in the training set have a disproportionately lower number of samples compared to others, which degrades classifier performance [25, 84, 105, 10, 15]. In security, commonly, the “attack” class has limited samples compared to the benign class, e.g., for malware detection, due to the challenges in sourcing attack data.
b) Under-represented in-class attack patterns (9/35 papers). Training dataset lacks enough instances of the different attack patterns that may arise [108, 47, 84, 51, 42]. Under-represented regions can exist within a class.
@@
c) Insufficient number of samples to fit a high-quality classifier (8/35 papers). Labeling or feature extraction processes are expensive in terms of computing effort, human effort (requires domain expertise), or other required resources. This is especially true in security where domain expertise is required to analyze software binary or run-time behavior to assign labels. This limits the amount of data collected to train an effective classifier [85, 51, 32, 98, 118].
@@
d) High dimensionality of features (3/35). As feature size increases, the number of samples needed for training increases exponentially, due to the ‘curse of dimensionality’ [12]. Such tasks tend to underperform, when the training data is limited [108, 111, 75].
@@
2) Concept drift (14/35 papers).   The classifier generalizes poorly to a test set distribution that has evolved and “drifted” away from the training distribution. Attackers are known to adapt over time to bypass defenses, leading to concept drift [51, 42, 111, 112, 10].
@@
3) Adversarial examples (6/35 papers).   Attackers exploit a core vulnerability of ML models by perturbing samples in the problem space (e.g., the malware binary to bypass classifiers) [20, 32, 19, 49, 15]. For example, Pierazzi et al. [78] identified problem-space transformations for Android malware to create evasive malware samples. Training data lack samples that capture such adversarial behavior, leading to degraded performance against adversarial inputs.
@@
In this work, we focus mainly on addressing the challenge of limited or biased training data, which is the most common challenge in our measurement study (Sections VII). In addition, we explore some preliminary directions to rapidly recover from concept drift (Section VIII). We leave the problem of adversarial examples for future work.
III Background and Related Work
III-A Problem setting and threat model
Security tasks and tabular data.   We study ML-based security classification tasks that use tabular data. Our analysis in Section II found that 28/35 articles utilize tabular data for ML-based security classification, highlighting its prevalence.
@@ III-A Problem setting and threat model
A tabular data sample can include $p$ continuous random variables: ${c_{1},...,c_{p}}$ and/or $q$ discrete (multinomial) random variables ${d_{1},...,d_{q}}$ that follow an unknown joint distribution. All training/validation/testing samples are encoded as tabular features. This is usually done by transforming the data from the problem space (original input space), such as a malware binary, into the feature space, e.g., a tabular feature, via static or dynamic analysis of the malware binary. We explore GenAI techniques that operate in a tabular feature space for training and generating data. Our focus is on supervised classifiers that need labeled datasets, covering binary and multi-class tasks. Tabular data can represent ‘attack’ and ‘benign’ samples, e.g., malware binary vs benign binary, or only attack samples, e.g., malware family attribution. Any ML model is applicable for classifier development.
@@
Defender and threat model.   The defender builds an ML-based security classifier to counter a cybersecurity threat. The task suffers from one or more of the data challenges discussed in Section II. The defender has access to a dataset to train and validate the performance of the security task. The existing training dataset is called the real dataset. Defender aims to use a GenAI tool to improve the performance of their security classifier before deployment and to maintain the performance after deployment, by updating the model over time. The GenAI tool is trained on the real data and then used to augment the real dataset with synthetic data to mitigate data challenges.
Prior work has mainly studied privacy-preserving synthetic data generation for ML tasks [93], to facilitate open data sharing while balancing privacy and utility.
However, privacy is not the focus of this study due to the substantial challenges in improving and sustaining classifier performance, with data sharing being a non-goal. We see it as complementary and a potential avenue for privacy-focused research.
III-B Research questions
@@
2) Can generating synthetic data using existing SOTA GenAI tools, and training a classifier on the augmented data boost the performance, compared to training only on real data?   We investigate the potential of the existing SOTA GenAI tools (designed for tabular data). We hypothesize that complementing real data with synthetic samples created using advanced GenAI tools can improve classification performance.
Note that existing GenAI tools are not tailor-made for security tasks, and no study has systematically tested these SOTA GenAI tools for security tasks. We also compare GenAI tools with traditional data augmentation strategies such as SMOTE.
@@
A reasonable question about our approach is how GenAI can help when training data is biased/limited. Our insight is that GenAI models that incorporate certain strategies can be trained on biased data to still produce effective synthetic samples. For example, we evaluate CTAB-GAN+ [117], a GAN-based GenAI scheme. CTAB-GAN+ uses strategies to prioritize more occurrence of rare samples (e.g., underrepresented attack patterns) in training set. It also uses special loss terms to weight rare samples during training. Conditional vectors are also used to guide the generator to generate synthetic samples for a specific class. Similarly, VAE-based GenAI models which use powerful priors over the latent space [94, 34, 56], can produce realistic variations of rare samples in the training set. That said, we do not expect all GenAI schemes to handle biased datasets equally well. Therefore, our work aims to provide a balanced perspective—both opportunities and challenges with using GenAI for security data augmentation.
@@
3) Can highly controlled synthetic data generation provide performance improvements for security tasks?   Existing GenAI tools provide limited control over the generation process—methods only allow generation conditioned on a class label. Such an uncontrolled synthesis that broadly mimics the real data is unlikely to capture complex biases within a class. We propose a novel GenAI scheme named Nimai that uses a VAE to facilitate highly controlled data generation. Conditioning on real samples, the tool provides greater control to correct biases and improve generalization.
@@
4) Can synthetic data from GenAI tools help with rapidly recovering from concept drift scenarios after deployment?   We leverage our controlled data synthesis tool, Nimai to enable fast recovery from concept drift scenarios. Given a limited number of labeled samples from the drifted distribution, Nimai can generate new synthetic samples that capture the drift. We study the efficacy of this rapid recovery.
III-C Related work

### v2

# Taming Data Challenges in ML-based Security Tasks Using Generative AI

@@ Abstract.
Machine learning-based supervised classifiers are widely used for security tasks, and their improvement has been largely focused on algorithmic advancements. Data challenges that negatively impact the performance of these classifiers have received limited attention. We address the following research question: Can developments in Generative AI (GenAI) address data challenges and improve classifier performance? We propose augmenting training datasets with synthetic data generated using GenAI techniques to improve classifier generalization. We evaluate this approach across 7 diverse security tasks using 6 state-of-the-art GenAI methods and introduce a novel GenAI scheme called Nimai that enables highly controlled data synthesis. We find that GenAI techniques can significantly improve the performance of security classifiers, achieving improvements of up to 32.6% even in severely data-constrained settings (only $\sim$180 training samples). Furthermore, we demonstrate that GenAI can facilitate rapid adaptation to concept drift post-deployment, requiring minimal labeling in the adjustment process. Despite successes, our study finds that some GenAI schemes struggle to initialize (train and produce data) on certain security tasks. We also identify characteristics of specific tasks, such as noisy labels, overlapping class distributions, and sparse feature vectors, which hinder performance boost using GenAI. We believe that our study will drive the development of future GenAI tools designed for security tasks.
generative AI, synthetic data, cybersecurity, security classification
journalyear: 2026copyright: ccconference: ACM Asia Conference on Computer and Communications Security; June 1–5, 2026; Bangalore, Indiabooktitle: ACM Asia Conference on Computer and Communications Security (ASIA CCS ’26), June 1–5, 2026, Bangalore, Indiadoi: 10.1145/3779208.3785264isbn: 979-8-4007-2356-8/26/06ccs: Security and privacy Systems securityccs: Computing methodologies Classification and regression trees
1. Introduction
@@
The application of machine learning (ML) to counteract cybersecurity threats is becoming increasingly prevalent. A widely adopted approach involves creating ML-based security classifiers through supervised learning to identify or categorize threats. These classifiers serve to identify or categorize malicious users or behaviors on online platforms (xu2021deep), malicious software (yang2021bodmas), network breaches (mirsky5kitsune), harmful network traffic/devices/entities (tekiner2022lightweight; cicids2017), software vulnerabilities (zhou2019devign), and web security hazards (yang2022wtagraph). These defense mechanisms emphasize lowering error rates (false alarms), enhancing generalization to unfamiliar test scenarios, and adding robustness against adaptive attackers. In pursuing these goals, primary attention has been paid to addressing the algorithmic challenges, which pertain to developing effective ML models and feature engineering techniques. However, the data challenges that affect the practical effectiveness of these defenses have received limited attention, resulting in significant limitations in further advancing performance. Our examination uncovers the following crucial data challenges in security (Section 2): significant class imbalance, insufficient representation of attack patterns, inadequate training samples, high-dimensional features, and concept drift (yang2021cade).
@@
In this study, we explore how recent advancements in AI can alleviate data-related challenges to enhance classification performance. This leads to our primary research inquiry: What are the challenges and opportunities associated with employing Generative Artificial Intelligence (GenAI) to address data challenges in machine learning (ML)-based security applications? GenAI models are capable of learning the distribution of a dataset and generating varied synthetic instances (sanh2019distilbert). A pivotal concept is utilizing carefully crafted synthetic data from GenAI to augment the existing training dataset of a security classifier, thereby enhancing generalization capabilities. Figure 1 (a) depicts the process of leveraging GenAI for data augmentation to refine performance before the classifier’s deployment. We specifically concentrate on tabular data tasks, as this represents a prevalent methodology for developing ML-based security classifiers. Examples of such tasks are presented in Sections 2 and 4. Employing GenAI to tackle data challenges in security is difficult for multiple reasons:
@@
  • Data synthesis in the security domain is inherently harder. Unlike other domains, e.g., computer vision (lu2022generative; ma2022comprehensive; chlap2021review), data in the security domain is inherently “adversarial” and captures behavior of an entity aiming to evade defenses. Thus, defenders typically have less knowledge of attack data distributions than of benign data. A GenAI scheme has to work with biased and limited attack data, as well as data with noisy samples (erroneous labels).
@@
  • Existing state-of-the-art GenAI models for tabular data are not tailor-made for security tasks. Several GenAI models for tabular data have emerged recently, e.g., TVAE (tvae/ctgan), CTAB-GAN+ (ctabgan+), TabDDPM (tabddpm), TabSyn (tabsyn), REaLTabFormer (solatorio2023realtabformer) and GReaT (GreaT). No study has systematically tested these GenAI schemes on security tasks. It is unclear how these GenAI schemes (for tabular data) can adapt to the data challenges in security.
@@
  • A GenAI scheme for security should not only address data challenges before deployment, but also after deployment. Since defenders and attackers are perpetually in an arms race, security classifiers are prone to drifting data or concept drift at test time (andresini2021insomnia). It is unclear how GenAI schemes can help with recovery from concept drift with minimal data labeling effort.
Figure 1. Overview of data augmentation using GenAI.:
@@
  • We start by highlighting key data challenges impacting development of security classifiers. We review 35 papers published in top security venues. We find 32/35 papers report at least one data challenge negatively impacting their classifier performance. We use our findings to select 7 security tasks suffering from these data challenges for evaluation. This includes diverse tasks: malware classification (yang2021bodmas; ahmadi2016novel; tekiner2022lightweight), OS fingerprinting (holland2021new), BGP hijacking detection (testart2019profiling), Tor-based malware traffic detection (dodia2022exposing), and web cookie privacy classification (bollinger2022automating).
  • We present the first systematic study of the effectiveness of SoTA GenAI schemes (designed for tabular data) to address data challenges and boost performance of security classifiers. We evaluate 6 SoTA GenAI schemes on our 7 security tasks. This includes GenAI models based on VAEs, LLMs, GANs and Diffusion models. We use synthetic data from GenAI schemes to augment training data for security tasks and boost performance.
  • We identify a missing capability in existing GenAI schemes (for tabular data)—none provide the control to target specific regions of the data manifold to alleviate complex biases. To fill this gap, we propose Nimai, a novel VAE-based scheme, which uses a discrete latent space to enable controlled sample-conditioned generations. Given an existing sample, Nimai can generate samples in its vicinity, providing more fine-grained control to the defender to address in-class bias issues. Our analysis shows that carefully crafted synthetic data improves classifier generalization by mitigating biases in the data.
@@
    Opportunities before deployment: This part focuses on mitigating data challenges before deployment of the classifier. We measure the performance gain (or degradation) from data augmentation compared to using only real training data. We obtain two key findings: (1) Existing GenAI schemes, among TVAE, CTAB-GAN+ and TabDDPM demonstrate performance improvement in 2 to 4 tasks, e.g., with up to 14.8% gain in the BGP hijacking detection task. (2) Our new scheme, Nimai which offers highly controlled data generation, achieved the best performance gain in 3/4 tasks. In an extremely data-challenged BGP hijacking detection task (<200<200 training samples), Nimai outperformed all baselines with up to a 32.6% gain. Our findings demonstrates the significant untapped potential of GenAI in security.
@@
    Challenges: Our analysis also revealed significant challenges with using GenAI: (1) We discover fundamental challenges with instantiating (training and generating data) existing GenAI models on our diverse tasks. 5/6 existing GenAI schemes failed to instantiate in at least one of the 7 tasks. For example, the LLM-based GenAI scheme called GReaT (GreaT) failed to generate data in 6/7 tasks due to scalability issues. (2) We identify 3 tasks with shared challenges of limited or biased data, and added complexities for GenAI: noisy labels, sparse features, and class overlap. Most GenAI schemes, including our new approach, fail to improve performance for these tasks, highlighting directions for future work.
We hope that our study will help to develop GenAI tools for security classification that address all identified challenges. Section 9 presents specific directions for future work. Our source code, datasets, and models are made available. [^1] https://github.com/secml-lab-vt/taming-data-challenges-security-ml.git
2. Data Challenges in Security Defenses
We conduct a study to characterize data challenges faced by researchers in ML-based cybersecurity tasks. We review 35 papers  [^2] Listed in Table 6 (Appendix). published in leading venues (e.g., IEEE S&P, USENIX Security and IMC) that developed at least one security classifier. The articles are semi-automatically selected based on their relevance to ML in security research. [^3] Our paper selection methodology is detailed in Appendix B. We find that 32/35 articles explicitly reported at least one data challenge that negatively affected classifier performance. We identify the following broad data challenges:
@@ 2. Data Challenges in Security Defenses
a) Imbalanced class distribution (10/35 papers). Certain classes in the training set have a disproportionately lower number of samples compared to others, which degrades classifier performance (cidon2019high; schuppen2018fanci; xu2019anatomy; barbero2022transcending; bollinger2022automating). In security, commonly, the “attack” class has limited samples compared to the benign class, e.g., for malware detection, due to the challenges in sourcing attack data. b) Under-represented in-class attack patterns (9/35 papers). Training dataset lacks enough instances of the different attack patterns that may arise (xu2021deep; ho2019detecting; schuppen2018fanci; jordaney2017transcend; han2021sigl). Under-represented regions can exist within a class.
@@
c) Insufficient number of samples to fit a high-quality classifier (8/35 papers). Labeling or feature extraction processes are expensive in terms of computing effort, human effort (requires domain expertise), or other required resources. This is especially true in security where domain expertise is required to analyze software binary or run-time behavior to assign labels. This limits the amount of data collected to train an effective classifier (banescu2017predicting; jordaney2017transcend; downing2021deepreflect; thirumuruganathan2022siraj; zhou2019devign).
@@
d) High dimensionality of features (3/35). As feature size increases, the number of samples needed for training increases exponentially, due to the ‘curse of dimensionality’ (bengio2000taking). Such tasks tend to underperform, when the training data is limited (xu2021deep; yang2021cade; park2019empirical).
@@
2) Concept drift (14/35 papers).   The classifier generalizes poorly to a test set distribution that has evolved and “drifted” away from the training distribution. Attackers are known to adapt over time to bypass defenses, leading to concept drift (jordaney2017transcend; han2021sigl; yang2021cade; yang2022wtagraph; barbero2022transcending).
@@
3) Adversarial examples (6/35 papers).   Attackers exploit a core vulnerability of ML models by perturbing samples in the problem space (e.g., the malware binary to bypass classifiers) (chen2020training; downing2021deepreflect; chen2021cost; jan2020throwing; bollinger2022automating). For example, Pierazzi et al. (pierazzi2020intriguing) identified problem-space transformations for Android malware to create evasive malware samples. Training data lack samples that capture such adversarial behavior, leading to degraded performance against adversarial inputs.
@@
In this work, we focus mainly on addressing the challenge of limited or biased training data, which is the most common challenge in our measurement study (Section 7). In addition, we explore preliminary directions to rapidly recover from concept drift (Section 8). We leave adversarial examples for future work.
3. Background and Related Work
3.1. Problem setting and threat model
@@
Security tasks and tabular data.   We study ML-based security classification tasks that use tabular data. Our analysis in Section 2 found that 28/35 articles utilize tabular data for ML-based security classification, highlighting its prevalence. A tabular data sample can include $p$ continuous random variables: $c_{1},...,c_{p}$} and/or $q$ discrete (multinomial) random variables $d_{1},...,d_{q}$} that follow an unknown joint distribution. All training/validation/testing samples are encoded as tabular features. This is usually done by transforming the data from the problem space (original input space), such as a malware binary, into the feature space, e.g., a tabular feature, via static or dynamic analysis of the malware binary. We explore GenAI techniques that operate in a tabular feature space for training and generating data. Our focus is on supervised classifiers that need labeled datasets, covering binary and multi-class tasks. Tabular data can represent ‘attack’ and ‘benign’ samples, e.g., malware binary vs benign binary, or only attack samples, e.g., malware family attribution. Any ML model is applicable for classifier development.
@@
Defender and threat model.   The defender builds an ML-based security classifier to counter a cybersecurity threat. The task suffers from one or more of the data challenges discussed in Section 2. The defender has access to a dataset to train and validate the performance of the security task. The existing training dataset is called the real dataset. Defender aims to use a GenAI tool to improve the performance of their security classifier before deployment and to maintain the performance after deployment, by updating the model over time. The GenAI tool is trained on the real data and then used to augment the real dataset with synthetic data to mitigate data challenges. The defender does not aim to re-design the security classifier, i.e., the focus is only on improving performance by addressing the training data challenges. We discuss related work that tightly integrates GenAI schemes into the design of the security classifier in Section 3.3, which is not our focus. Prior work has also studied privacy-preserving synthetic data generation for ML tasks (stadler2022synthetic), to facilitate open data sharing while balancing privacy and utility. However, privacy is not the focus of this study due to the substantial challenges in improving and sustaining classifier performance, with data sharing being a non-goal. We see it as complementary and a potential avenue for privacy-focused research.
3.2. Research questions
@@
2) Can generating synthetic data using existing SOTA GenAI tools, and training a classifier on the augmented data boost the performance, compared to training only on real data?   We investigate the potential of the existing SOTA GenAI tools (designed for tabular data). We hypothesize that complementing real data with synthetic samples created using advanced GenAI tools can improve classification performance. Note that existing GenAI tools are not tailor-made for security tasks, and no study has systematically tested these SOTA GenAI tools for security tasks. We also compare GenAI tools with traditional data augmentation strategies such as SMOTE.
@@
A reasonable question about our approach is how GenAI can help when training data is biased/limited. Our insight is that GenAI models that incorporate certain strategies can be trained on biased data to still produce effective synthetic samples. For example, we evaluate CTAB-GAN+ (ctabgan+), a GAN-based GenAI scheme. CTAB-GAN+ uses strategies to prioritize more occurrence of rare samples (e.g., underrepresented attack patterns) in training set. It also uses special loss terms to weight rare samples during training. Conditional vectors are also used to guide the generator to generate synthetic samples for a specific class. Similarly, VAE-based GenAI models which use powerful priors over the latent space (stocksieker2024data; fajardo2018vos; kim2024tvariational), can produce realistic variations of rare samples in the training set. That said, we do not expect all GenAI schemes to handle biased datasets equally well. Therefore, our work aims to provide a balanced perspective—both opportunities and challenges with using GenAI for security data augmentation.
@@
3) Can highly controlled synthetic data generation provide performance improvements for security tasks?   Existing Gen- AI tools provide limited control over the generation process— methods only allow generation conditioned on a class label. Such an uncontrolled synthesis that broadly mimics the real data is unlikely to capture complex biases within a class. We propose a novel GenAI scheme named Nimai that uses a VAE to facilitate highly controlled data generation. Conditioning on real samples, the tool provides greater control to correct biases and improve generalization.
@@
4) Can synthetic data from GenAI tools help with rapidly recovering from concept drift scenarios after deployment?   We leverage our controlled data synthesis tool, Nimai to enable fast recovery from concept drift scenarios. Given a limited number of labeled samples from the drifted distribution, Nimai can generate new synthetic samples that capture the drift.
3.3. Related work